0. If the height (in inches) of newborn babies has the N (18, 1) distribution, what is the probability that the height of a newborn baby is between 17 and 20 inches? this distribution, what is the probability that the height of a newborn baby is n neonle suf- What is the distribution of height in centimeters ( inch 2.54 em)? Using between 43.18 cm (17 inches) and 50.80 cm (20 inches)?
Answer #1

10. P(17<x<20)=P(\frac{17-18}{1}<z<\frac{20-18}{1})=P(-1<z<2)=0.8186

Now converting inches to centimeters

x\sim N(45.72,2.54^2)

So P(43.18<x<50.80)=P(\frac{43.18-45.72}{2.54}<z<\frac{50.80-45.72}{2.54})=P(-1<z<2)=0.8186